# Numerical boundary flux functions that give provable bounds for nonlinear initial boundary value problems with open boundaries

### v1

@@ 5.2 A derivation using the general theory
The generic statement (4.6) yields the same numerical boundary flux, provided we can form the (nonlinear) characteristic variables $\mathbf{W}$ by finding the matrix $\underline{\textnormal{T}}$, which then guides the ansatz for boundary data $\mathbf{G}$. As before, we look only at the $u>0$ case. At the left boundary for the scalar equation, $\underline{\textnormal{I}}^{-}\leftarrow 1$. For the Burgers equation, $\mathbf{U}=\mathbf{V}\leftarrow u$, i.e., the conservative and entropy variables are the same, so the transformation $\underline{\textnormal{T}}\leftarrow 1$. Furthermore, the boundary data according to the continuous analysis (5.4) is
@@
Thus, the formalism of (4.6) leads to a numerical boundary flux function (5.17) that is identical to the one derived from first principles in (5.11). Note that the creation of the boundary flux function from (4.6) requires knowledge of the characteristic variables $\mathbf{W}$, the wave speeds $\underline{\textnormal{$\Lambda$}}$, and the boundary terms $\mathbf{G}$.
@@ 6 Stable nonlinear boundary fluxes for the shallow water equations
where $h$ is the water height, $g$ is the gravitational constant, and $v_{1},v_{2}$ are the velocities in the $x$ and $y$ directions [36]. The normal flux at the boundary $\partial\Omega$ with normal vector $\accentset{\,\rightarrow}{n}=(n_{1},n_{2})^{T}$ is
@@ 7 Numerical results
We apply the newly derived boundary fluxes to compute solutions of the Burgers and shallow water equations with DGSEM approximations. The spatial discretizations for the split form DGSEM are available in Trixi.jl [43, 2] and TrixiShallowWater.jl [44]. For time integration we use CFL-based time stepping with the five-stage, four-order explicit Runge-Kutta method of Carpenter and Kennedy [45] implemented in OrdinaryDiffEq.jl [46]. The unstructured curvilinear quadrilateral meshes were constructed with HOHQMesh [47, 48]. We use Plots.jl [49] and ParaView [50] to visualize the results. All source code needed to reproduce the numerical experiments is available online in our reproducibility repository [51].
@@ 7.1 Burgers’ equation
We take the final time to be $t=120$, $\text{CFL}=0.75$, and use one of three boundary fluxes to impose the boundary condition: (1) The entropy conservative (EC) flux [30], (2) The LLF flux [35], and (3) the new boundary flux from (5.11). The computation at all interior element interfaces uses the entropy conservative flux so the approximation is dissipation-free in the interior, and the only dissipation is introduced at the left and right physical boundaries.
For each test we use a five element mesh with polynomials of degree $N=7$ on each element. The EC flux is not an upwind Riemann solver and does not consider the characteristics in its design. This, unsurprisingly, makes it an especially poor choice for open boundary conditions as there is no bound, cf. Rem. 5.13. We see in Fig. 1(a) that the solution computed with the EC flux at the open boundaries exhibits instability. Imposing the inflow-outflow boundary conditions with the LLF flux does run successfully to the final time for this problem, see Fig. 1(b) even though the inflow condition is not provably bounded, cf. Rem. 5.13. (There is outflow dissipation however to mitigate that fact.) The boundary flux (5.11) produces Fig. 1(c), which looks identical to the LLF result. At the final time the $L_{2}$ error with the new flux result is $8.80419344\cdot 10^{-7}$ and the LLF result has an $L_{2}$ error of $8.80425611\cdot 10^{-7}$, which are the same to four significant digits.
@@
Figure 1: The solution (7.1) of Burgers’ equation at final time $t=120$ with polynomials of degree seven on a five element mesh.
@@ 8 Concluding remarks
We have shown how to interpret nonlinearly stable open boundary treatments as numerical boundary flux functions. These fluxes penalize (nonlinear) characteristic-type variables at inflow-outflow boundaries. Incorporated into a high-order split form DGSEM, the nonlinear solution including open boundaries is provably bounded by specified, external data.

### v2

@@ 5.2 A derivation using the general theory
The generic statement (4.6) yields the same numerical boundary flux, provided we can form the (nonlinear) characteristic variables $\mathbf{W}$ by finding the matrix $\underline{\textnormal{T}}$, which then guides the ansatz for boundary data function $\mathbf{G}$. As before, we look only at the $u>0$ case. At the left boundary for the scalar equation, $\underline{\textnormal{I}}^{-}\leftarrow 1$. For the scalar Burgers equation, $\mathbf{U}=\mathbf{V}\leftarrow u$, i.e., the conservative and entropy variables are the same, so the transformation $\underline{\textnormal{T}}\leftarrow 1$. Furthermore, the boundary data according to the continuous analysis (5.4) is
@@
Thus, the formalism of (4.6) leads to a numerical boundary flux function (5.17) that is identical to the one derived from the energy method in (5.11). Note that the creation of the boundary flux function from (4.6) requires knowledge of the characteristic variables $\mathbf{W}$, the wave speeds $\underline{\textnormal{$\Lambda$}}$, and the boundary terms $\mathbf{G}$.
Remark 5.3.
The final form of the numerical flux derived via the energy method (5.11) or from the generic result of Thm. 4.2 in (5.17) is a function of the external solution state $u_{\textrm{ext}}$. It does not explicitly contain the boundary data function $\mathbf{G}$, but is derived from it, as shown in (5.4) and (5.14).
@@ 6 Stable nonlinear boundary fluxes for the shallow water equations
where $h$ is the water height, $g$ is now the gravitational constant, and $v_{1},v_{2}$ are the velocities in the $x$ and $y$ directions [39]. The normal flux at the boundary $\partial\Omega$ with outward normal vector $\accentset{\,\rightarrow}{n}=(n_{1},n_{2})^{T}$ is
@@ 6.2 Nonlinear numerical boundary flux functions for the shallow water equations
Again, the numerical boundary fluxes $\mathbf{F}^{*}_{n}$ are functions of the internal and external solution states and will not explicitly contain the boundary term $\mathbf{G}$, see Remark 5.3.
@@ 7 Numerical results
We apply the newly derived boundary fluxes to compute solutions of the Burgers and shallow water equations with DGSEM approximations. The spatial discretizations for the split form DGSEM are available in Trixi.jl [45, 2] and TrixiShallowWater.jl [46]. For time integration we use CFL-based time stepping with the five-stage, four-order explicit Runge-Kutta method of Carpenter and Kennedy [47] implemented in OrdinaryDiffEq.jl [48]. The unstructured curvilinear quadrilateral meshes were constructed with HOHQMesh [49, 50]. We use Plots.jl [51] and ParaView [52] to visualize the results. The implementations of the nonlinearly stable boundary fluxes derived here as well as all source code needed to reproduce the numerical experiments is available online in our reproducibility repository [53].
@@ 7.1 Burgers’ equation
We take the final time to be $t=120$, $\text{CFL}=0.75$, and use one of three boundary fluxes to impose the boundary condition: (1) The entropy conservative (EC) flux [32], (2) The LLF flux [38], and (3) the new boundary flux from (5.11). The computation at all interior element interfaces uses the entropy conservative (EC) flux so the approximation is dissipation-free in the interior, and the only dissipation is introduced at the left and right physical boundaries.
For each test we use a five element mesh with polynomials of degree $N=7$ on each element. The EC flux is not an upwind Riemann solver and does not consider the characteristics in its design. This, unsurprisingly, makes it an especially poor choice for open boundary conditions as there is no bound, cf. Rem. 5.13. We see in Fig. 1(a) that the solution computed with the EC flux at the open boundaries exhibits instability. Imposing the inflow-outflow boundary conditions with the LLF flux does run successfully to the final time for this problem, see Fig. 1(b) even though the inflow condition is not provably bounded, cf. Rem. 5.13. (There is outflow dissipation, however, to mitigate that fact.) The boundary flux (5.11) produces Fig. 1(c), which looks identical to the LLF result. At the final time, the $L_{2}$ error with the new flux result is $8.80419344\cdot 10^{-7}$ and the LLF result has an $L_{2}$ error of $8.80425611\cdot 10^{-7}$, which are the same to four significant digits.
@@
Figure 1: The solution (7.1) of Burgers’ equation at final time $t=120$ with polynomials of degree seven on a five element mesh. All element interfaces use the EC flux; so the approximation is dissipation-free in the interior. Thus, the only dissipation in the approximation is introduced at the left and right physical boundaries by the selected boundary fluxes.
@@ 8 Concluding remarks
We have shown how to interpret nonlinearly stable open boundary treatments as numerical boundary flux functions. These fluxes penalize (nonlinear) characteristic-type variables at inflow-outflow boundaries. Incorporated into a high-order split form DGSEM, the nonlinear solution including open boundaries is provably bounded by specified external data.